A financial planner recommended that part of the client’s 600000 Pesos cash account be invested in preferred stock earning 9% annual simple interest. The remainder of the client’s cash was placed in Treasury bong earning 7% annual simple interest. The total annual interest earned from two investments was 51000 Pesos. Find the amount invested at each rate.
Answer by stanbon(75887)

A financial planner recommended that part of the client’s 600000 Pesos cash account be invested in preferred stock earning 9% annual simple interest. The remainder of the client’s cash was placed in Treasury bong earning 7% annual simple interest. The total annual interest earned from two investments was 51000 Pesos. Find the amount invested at each rate.
-------
p + t = 600000
9p + 7t = 5100000
-----------------------
Modify for elimination.
9p + 9t = 5400000
9p + 7t = 5100000
-------------------------
2t = 300000
t = $150,000
---
p = $450,000
-------------
Cheers,
Stan H.
----------